The following trial balance pertains to Benji’s Grocery as of January 1, 2018:

Account Title Beginning Balances
Cash $ 64,000
Accounts receivable 12,000
Merchandise inventory 90,000
Accounts payable 7,500
Common stock 89,000
Retained earnings 69,500

The following events occurred in 2018. Assume that Benji’s uses the periodic inventory method.

  1. Purchased land for $30,000 cash.

  2. Purchased merchandise on account for $230,000, terms 1/10, n/45.

  3. Paid freight of $2,100 cash on merchandise purchased FOB shipping point.

  4. Returned $8,600 of defective merchandise purchased in Event 2.

  5. Sold merchandise for $186,000 cash.

  6. Sold merchandise on account for $236,000, terms 2/10, n/30.

  7. Paid cash within the discount period on accounts payable due on merchandise purchased in Event 2. *Hint: there are two separate journal entries for this one, 7A and 7B.

  8. Paid $28,500 cash for selling expenses.

  9. Collected $156,000 of the accounts receivable from Event 6 within the discount period.

  10. Collected $56,000 of the accounts receivable but not within the discount period.

  11. Paid $17,100 of other operating expenses.

  12. A physical count indicated that $48,300 of inventory was on hand at the end of the accounting period.


Required:

  1. Record the preceding transactions in a horizontal statements model. In the Cash Flows column, use OA to designate operating activity, IA for investing activity, FA for financing activity, NC for net change in cash and NA to indicate accounts not affected by the event. The beginning balances have been recorded as an example.
Answer #1

Ans- Record the transactions in the horizontal statements model as follows:-

Balance Sheet Income Statement
Event No. Assets = Liabilities
Cash + Accounts Receivable + Merchandise Inventory + Land = Accounts Payable + Common Stock + Retained Earnings Revenue - Expenses = Net Income Statement of Cash Flows
Beginning Balance 64,000 + 12,000 + 90,000 + = 7,500 + 89,000 + 69,500 NA - NA = NA NA NA
1 (30,000) + + + 30,000 = + + - = (30,000) IA
2 + + 230,000 + = 230,000 + + - = NA
3 (2,100) + + 2,100 + = + + - = (2,100) OA
4 + + (8,600) + = (8,600) + + - = NA
5 186,000 + + + = + + 186,000 186,000 - = 186,000 186,000 OA
6 + 236,000 + + = + + 236,000 236,000 - = 236,000 NA
7a (219,186) + + + = (221,400) + + - = (219,186) OA
7b. + + (2,214) + = + + - = NA
8 (28,500) + + + = + + (28,500) - 28,500 = (28,500) (28,500) OA
9 152,880 + (156,000) + + = + + (3,120) - 3,120 = (3,120) 152,880 OA
10 56,000 + (56,000) + + = + + - = 56,000 OA
11 (17,100) + + + = + + (17,100) - 17,100 = (17,100) (17,100) OA
12 + + 48,300 + = + + 48,300 - =
Total 161,994 + 36,000 + 359,586 + 30,000 = 7,500 + 89,000 + 491,080 422,000 - 48,720 = 373,280